OpenStudy (anonymous):

Suppose the circumference of a circle is 4π. What is its area?

OpenStudy (igreen):

\(\sf C = 2 \pi r\) Find the radius first, then plug it into the formula to find the area.

OpenStudy (igreen):

Plug in what we know: \(\sf 4 \pi = 2 \pi r\) Divide 2 pi to both sides

OpenStudy (16shuston):

What is the formula for the circumference? When you know this you have to solve for r then take the formula for the area of the circle and plug in r

OpenStudy (anonymous):

\[\frac{ 4 \pi }{ 2 \pi } = \frac{ 2 \pi r }{ 2 \pi } \]

OpenStudy (anonymous):

@iGreen does r=2

OpenStudy (igreen):

Yes!

OpenStudy (anonymous):

yay!

OpenStudy (igreen):

Formula for the area of a circle: \(\sf A = \pi r^2\) Plug in 2 for the radius: \(\sf A = \pi (2^2)\)

OpenStudy (igreen):

2^2 = ?

OpenStudy (anonymous):

\[A=\pi(2^{2})=4 \times3.14= 6.28??\]

OpenStudy (igreen):

Yes, but I don't think they want you to simplify pi..

OpenStudy (igreen):

If they do, 6.28 is correct..but if they don't, it's \(\sf 4 \pi\).

OpenStudy (anonymous):

ok i'll let know in a sec

OpenStudy (anonymous):

ya its \[4 \pi\]

OpenStudy (anonymous):

Thx so much i get this now i wrote down notes to understand @iGreen

OpenStudy (igreen):

Yay! \(\Huge\ddot\smile\)